# Calibrating Deep Neural Network using Euclidean Distance

### v1

@@ 5. Experimental Results
We evaluate the effectiveness of our proposed method on both image and text classification tasks. For image classification, we utilize datasets such as CIFAR-10/100 (Krizhevsky, 2009), Tiny-ImageNet (Deng et al., 2009), and CheXNet (Rajpurkar et al., 2017) to evaluate calibration of models including ResNet50 (He et al., 2016a), ResNet-100 (He et al., 2016b), Wide-ResNet-26-10 (Zagoruyko and Komodakis, 2016), and DenseNet-121 (Huang et al., 2017). For text classification, we employ the 20 Newsgroups dataset (Lang, 1995), web news dataset AG News (Zhang et al., 2015) and FinSen (Liang et al., 2024), training with global-pooling CNN (Lin et al., 2014). More, we include Out-of-Distribution (OoD) datasets to evaluate the robustness of the models: the SVHN dataset (Netzer et al., 2011), which consists of street view house numbers, and the CIFAR-10-C dataset (Hendrycks and Dietterich, 2019), a corrupted version of the CIFAR-10.
@@
Experiment Setup Our experiment setup aligns with prior work and public code by Mukhoti et al. (Mukhoti et al., 2020a) implemented using PyTorch. CIFAR-10/100: Trained ResNet-50, ResNet-110, WideResNet, DenseNet for 350 epochs, with a learning rate set to 0.1 for the first 150 epochs, 0.01 for the following 100 epochs, and 0.001 for the remaining epochs. Tiny-ImageNet: Trained ResNet for 100 epochs, with a learning rate set to 0.1 for the first 40 epochs, 0.01 for the following 20 epochs, and 0.001 for the remaining epochs. 20 Newsgroups: Trained Global Pooling CNN for 50 epochs, with a learning rate set to 0.001, beta set from 0.9 to 0.999. We used Glove word embedding (glove.6B.100d) to train and evaluated by the best validated models. CheXNet: Trained DenseNet for 100 epochs with a ReduceLROnPlateau scheduler (patience = 5). For optimization, we used SGD with a weight decay of $5\times 10^{-4}$ and momentum of 0.9 for CIFAR-10/100 and Tiny-ImageNet, Adam optimizer with a learning rate of 0.001 for 20Newsgroup, and Adam optimizer with a learning rate of 0.0001 for CheXNet with $1\times 10^{-5}$ weight decay. The batch sizes are all set to 128 but 32 for CheXNet. All experiments were conducted on NVIDIA GeForce RTX 3060 Ti and NVIDIA GeForce RTX 4090, with random seeds set to 1. For temperature scaling results, the temperature parameter $T$ was optimized through a grid search over the range $T\in[0,0.1,0.2,\ldots,10]$ on the validation set, based on the best post-temperature-scaling ECE. The exact temperature parameter was also used for other metrics, such as AdaECE, Classwise-ECE and smCE.
5.1. Calibration Performance
5.1.1. Calibration Error
@@
Table 1. $\downarrow$ ECE (%) for various methods both before (pre) and after (post) applying temperature scaling (Bin = 15)  [^3] The smCE metrics is based on the approach presented by Błasiok and Nakkiran (Błasiok and Nakkiran, 2023), with the corresponding implementation publicly available on GitHub (Błasiok, Jarosław and Nakkiran, Preetum, 2024)..
Dataset Model Weight Decay(Guo et al., 2017) Brier Loss(Brier, 1950) MMCE(Kumar et al., 2018) Label Smooth(Szegedy et al., 2016) Focal Loss - 53(Mukhoti et al., 2020a) Dual Focal(Tao et al., 2023) Focal Calibration
@@
Table 2. $\downarrow$ Classification error (%) on test set across different methods.
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal(Tao et al., 2023) Focal Calibration
@@
Table 3. $\downarrow$ smCE (%) for various methods both before (pre) and after (post) applying temperature scaling 55footnotemark: 5.
Dataset Model Weight Decay(Guo et al., 2017) Brier Loss(Brier, 1950) MMCE(Kumar et al., 2018) Label Smooth(Szegedy et al., 2016) Focal Loss - 53(Mukhoti et al., 2020a) Dual Focal(Tao et al., 2023) Focal Calibration
@@ 5.1.2. Reliability Diagram
The Fig. 9 for ResNet-110 trained on CIFAR-10 show the reliability diagram for different loss functions: Cross Entropy, FLSD-53, Dual Focal, and Focal Calibration. The Cross Entropy loss shows a significant gap between the expected and actual accuracy, especially in the higher confidence bins indicated poor calibration. Focal Calibration Loss achieves the best calibration among the methods, with the lowest ECE and a relatively low MCE. The expected and actual accuracies are closely aligned, indicating excellent calibration. Additionally, it has the lowest test error, demonstrating that FCL not only improves calibration but also maintains or improves classification performance. The Fig. 10 for ResNet-50 trained on CIFAR-10 with epochs range from 100 to 300 also proves it.
5.2. Classification Performance
5.2.1. Classification Error
@@
The Tab. 2 presents the classification error (%) on the test set for various methods across different datasets and models. FCL consistently achieves the lowest classification error in most cases across all datasets and models, demonstrating its effectiveness in reducing test errors.
5.3. Gradient-Based Localization
5.3.1. Grad-CAM Heatmaps on CheXNet
@@
We evaluate the effects of Cross Entropy, Focal Loss, and Focal Calibration Loss on localization performance using Grad-CAM heatmaps, a popular visualization technique for identifying regions of interest in deep learning models (Selvaraju et al., 2017). Fig. 1 presents a comparison of heatmaps generated with each loss function applied to CheXNet (Rajpurkar et al., 2017), a deep learning model for chest X-ray classification. For Cross Entropy, the heatmaps show broader regions of activation, which suggest that the model has less certainty in localizing the anomalies. In contrast, focal loss demonstrates more concentrated regions of activation. This suggests that focal loss sharpens the model’s focus on challenging examples. However, while the activations are more focused, they may not always be aligned with precise boundaries, which could affect the confidence in anomaly localization. Focal Calibration Loss, on the other hand, provides the most well-defined and precise activation areas. The heatmaps generated show a balanced focus, particularly in difficult cases, while also improving probability calibration. This results in clearer boundaries and more accurate localization of anomalies lead to more confident and reliable predictions, making it especially suited for medical diagnosis tasks where precise localization and confidence in predictions are critical. Thus, FCL not only enhances model calibration but also refines anomaly localization, offering substantial benefits for applications in medical imaging where accurate and confident diagnoses are of utmost importance.
5.4. Robustness on OoD Data Shift
@@
We assess the robustness of models trained on in-distribution CIFAR-10 and evaluated on two distinct out-of-distribution (OoD) datasets: SVHN and CIFAR-10-C. Tab. 4 presents the performance of Area Under the Receiver Operating Characteristic curve (AUROC) under data shifts. Across most datasets, FCL consistently achieves the highest AUROC, outperforming other models in handling OoD data. This indicates that FCL enhances the model’s ability to distinguish between in-distribution and out-of-distribution samples effectively. This suggests that the robustness improvements introduced by FCL generalize well across different types of OoD shifts.
@@
Table 4. Robustness on OoD Dataset Shift. $\uparrow$ AUROC (%) for shifting CIFAR-10 (in-distribution) to SVHN and CIFAR-10-C (OoD).
@@
Dataset Model Weight Decay(Guo et al., 2017) Brier Loss(Brier, 1950) MMCE(Kumar et al., 2018) Label Smooth(Szegedy et al., 2016) Focal Loss - 53(Mukhoti et al., 2020a) Dual Focal(Tao et al., 2023) Focal Calibration
@@
5.5. Discussion
5.5.1. Faster convergence
@@
Figure 4. Normalized test loss across training epochs for different loss functions on CIFAR-10 using ResNet-110. The comparison includes Cross Entropy, Focal Loss ($\gamma$=3.0), Dual Focal Loss ( $\gamma$=2.0), and Focal Calibration Loss ($\gamma$=3.0). Focal Calibration Loss with faster convergence and stability.
[…]
tbc
5.5.2. Balance between $\gamma$ and $\lambda$
@@
We compared the FCL and FL with different $\lambda$ values while keeping $\gamma=3.0$ constant by metrics including NLL, ECE, AdaECE, and Classwise-ECE both before and after temperature scaling as well as test error for each $\lambda$ in Fig. 7. NLL decreases initially as $\lambda$ increases from 0 to 0.5 but shows fluctuations and a slight increase beyond $\lambda=1.5$. For post-temperature scaling, NLL values are generally lower, indicating improved calibration. ECE decreases significantly when $\lambda$ is 0.5, indicating improved calibration. However, calibration error increases for higher $\lambda$ values, suggesting that too much regularization might degrade calibration performance as depicted in Fig. 7. The experiment shows that $\lambda=0.5$ provides the best calibration performance across all ECE metrics before and after temperature scaling. Adjusting $\gamma$ influences how much the model concentrates on hard examples—higher $\gamma$ increases this focus, which can improve calibration but may also lead to overfitting if excessive. Modifying $\lambda$ controls the weight of the calibration loss component; increasing $\lambda$ emphasizes probability calibration by penalizing deviations between predicted and true probabilities, but overly large values can reduce classification accuracy. Therefore, appropriate tuning of these hyperparameters is crucial to achieving a desirable balance between calibration and classification effectiveness. We provide table of $\gamma$ and $\lambda$ selection with best calibration error performance in Appendix A.10.
5.5.3. Temperature scaling and post-processing gap
@@
To empirically validate the theoretical reduction in post-processing gap as Theorem 8, we compare the ECE/smCE/AdaECE/Classwise-ECE of models trained with focal loss and FCL before and after temperature scaling. As shown in Tab. 3, Tab. 5, Tab. 6 and Tab.7, the focal loss model exhibits a significant reduction in ECE after temperature scaling, indicating a substantial post-processing gap. In contrast, the FCL model shows minimal change in ECE after temperature scaling closing with 1.0, particularly in smCE, confirming that FCL effectively reduces the post-processing gap during training.
@@
Figure 5. Different hyperparameters trained on CIFAR-10 with ResNet-110. The plots show the performance of the FCL with varying calibration parameter $\lambda$ ($\gamma=3.0$) across different metrics: NLL, ECE, AdaECE, and Classwise-ECE. Each plot includes before (blue solid lines) and after (orange dashed lines) temperature scaling, alongside the Test Error (red dotted lines) 77footnotemark: 7.
[…]
tbc
6. Related Work
@@
Improving the calibration of deep neural networks (DNNs) is critical for reliable uncertainty estimation. A common strategy involves post-hoc calibration methods, which adjust a pre-trained model’s output probabilities without altering its architecture or training procedure. Platt scaling (Platt et al., 1999), originally designed for binary classifiers, was extended to multi-class settings by applying a linear transformation to logits to improve calibration. More recent methods like Temperature Scaling (Guo et al., 2017) introduce a temperature parameter to adjust the predicted logits, effectively reducing the Expected Calibration Error (ECE) across various models. Histogram Binning (Zadrozny and Elkan, 2002) discretizes the probability space into bins to align confidence scores with empirical frequencies. While effective, these methods require held-out validation data and assume the test distribution matches the calibration data.
6.1. Calibration Strategies
@@
Regularization-based approaches aim to improve calibration during training. Label Smoothing (Müller et al., 2019; Szegedy et al., 2016) mitigates overconfidence by blending true labels with a uniform distribution, softening hard targets to enhance generalization and uncertainty estimation. Additionally, auxiliary losses like the Maximum Mean Calibration Error (MMCE) (Kumar et al., 2018) explicitly minimize the difference between predicted and actual confidence levels. Bayesian techniques offer a principled way to quantify uncertainty. For instance, Deep Ensembles (Lakshminarayanan et al., 2017) aggregate predictions from multiple independently trained models to yield well-calibrated confidence estimates, though they are computationally intensive and less practical in resource-limited settings. Focal Loss (Lin et al., 2017), designed to handle class imbalance in object detection, also affects calibration by concentrating on hard-to-classify examples, which can reduce overconfidence but sometimes result in underconfident predictions (Charoenphakdee et al., 2021). Dual Focal Loss (DFL) (Tao et al., 2023) improves upon this by maximizing the margin between the true class logit and the next highest logit, balancing overconfidence and underconfidence to enhance calibration. Recently, smoothing-based methods have been introduced to provide continuous calibration measures. The Smooth Calibration Error (smCE) (Błasiok and Nakkiran, 2023) uses kernel-based smoothing to create a differentiable measure of calibration error, overcoming the limitations of discrete binning in traditional ECE calculations.
6.2. Calibration in Transformers and LLMs
Calibration of transformer-based language models has been evaluated across various natural language processing tasks, including machine translation (Kumar and Sarawagi, 2019), question answering (Jiang et al., 2020), and selective prediction (Varshney et al., 2022; Lu et al., 2022). Recently, with the growing prominence of large-scale generative language models (LLMs), studies have begun to examine their calibration properties (Lin et al., 2022; Kuhn et al., 2023; Tian et al., 2023; Xiong et al., 2023; Huang et al., 2024) and zero-shot/few-shot learning (Brown, 2020). To enhance text generation quality and improve models’ handling of ambiguity, more studies have focused on addressing miscalibration of token-level probabilities by implementing calibration techniques that enable models to express uncertainty more accurately (Kumar and Sarawagi, 2019; Mielke et al., 2022; Lin et al., 2022; Xu et al., 2023; Zhou et al., 2023).
7. Conclusions

### v2

@@ 5 Experimental Results
We evaluate the effectiveness of our proposed method on both image and text classification tasks. For image classification, we utilize datasets such as CIFAR-10/100 (Krizhevsky, 2009), Tiny-ImageNet (Deng et al., 2009), and CheXNet (Rajpurkar et al., 2017) to evaluate calibration of models including ResNet50 (He et al., 2016a), ResNet-100 (He et al., 2016b), Wide-ResNet-26-10 (Zagoruyko & Komodakis, 2016), and DenseNet-121 (Huang et al., 2017). For text classification, we employ the 20 Newsgroups dataset (Lang, 1995), web news dataset AG News (Zhang et al., 2015) and FinSen (Liang et al., 2024), training with global-pooling CNN (Lin et al., 2014). More, we include Out-of-Distribution (OoD) datasets to evaluate the robustness of the models: the SVHN dataset (Netzer et al., 2011), which consists of street view house numbers, and the CIFAR-10-C dataset (Hendrycks & Dietterich, 2019), a corrupted version of the CIFAR-10.
@@
Experiment Setup Our experiment setup aligns with prior work and public code by Mukhoti et al. (Mukhoti et al., 2020a) implemented using PyTorch. CIFAR-10/100: Trained ResNet-50, ResNet-110, WideResNet, DenseNet for 350 epochs, with a learning rate set to 0.1 for the first 150 epochs, 0.01 for the following 100 epochs, and 0.001 for the remaining epochs. Tiny-ImageNet: Trained ResNet for 100 epochs, with a learning rate set to 0.1 for the first 40 epochs, 0.01 for the following 20 epochs, and 0.001 for the remaining epochs. 20 Newsgroups: Trained Global Pooling CNN for 50 epochs, with a learning rate set to 0.001, beta set from 0.9 to 0.999. We used Glove word embedding (glove.6B.100d) to train and evaluated by the best validated models. CheXNet: Trained DenseNet for 100 epochs with a ReduceLROnPlateau scheduler (patience = 5). For optimization, we used SGD with a weight decay of $5\times 10^{-4}$ and momentum of 0.9 for CIFAR-10/100 and Tiny-ImageNet, Adam optimizer with a learning rate of 0.001 for 20Newsgroup, and Adam optimizer with a learning rate of 0.0001 for CheXNet with $1\times 10^{-5}$ weight decay. The batch sizes are all set to 128 but 32 for CheXNet. All experiments were conducted on NVIDIA GeForce RTX 3060 Ti and NVIDIA GeForce RTX 4090, with random seeds set to 1. For temperature scaling results, the temperature parameter $T$ was optimized through a grid search over the range $T\in[0,0.1,0.2,\ldots,10]$ on the validation set, based on the best post-temperature-scaling ECE. The exact temperature parameter was also used for other calibration metrics.
5.1 Calibration Performance
5.1.1 Calibration Error
@@
Table 1: $\downarrow$ ECE (%) for various methods both before (pre) and after (post) applying temperature scaling (Bin = 15)  [^4] We follow the experimental methodology of (Mukhoti et al., 2020a), utilizing their publicly available code (Mukhoti et al., 2020b). For the Dual Focal Loss, we initially set $\gamma=5$ based on preliminary tuning reported by (Tao et al., 2023) to ensure a fair comparison. Since specific $\gamma$ values for other models and datasets were not provided, we performed cross-validation to determine the optimal $\gamma$ for each experiment. These optimized $\gamma$ values may influence the observed results by enhancing model calibration and performance. Optimal temperature scaling values determined by validation set (included in brackets) close to 1.0 indicate well-calibrated predictions requiring minimal adjustment..
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
@@
Table 2: $\downarrow$ Classification error (%) on test set across different methods.
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
@@
Table 3: $\downarrow$ smCE (%) for various methods both before (pre) and after (post) applying temperature scaling  [^5] The smCE metrics is based on the approach presented by (Błasiok & Nakkiran, 2023), with implementation publicly on GitHub (Błasiok, Jarosław and Nakkiran, Preetum, 2024)..
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
@@ 5.1.2 Reliability Diagram
The Fig. 9 for ResNet-110 trained on CIFAR-10 show the reliability diagram for different loss functions: Cross Entropy, FLSD-53, Dual Focal, and Focal Calibration. The Cross Entropy loss shows a significant gap between the expected and actual accuracy, especially in the higher confidence bins indicated poor calibration. Focal Calibration Loss achieves the best calibration among the methods, with the lowest ECE and a relatively low MCE. The expected and actual accuracies are closely aligned, indicating excellent calibration. Additionally, the lowest test error is demonstrating that FCL not only improves calibration but also maintains or improves classification performance as Fig. 8 for ResNet-50 trained on CIFAR-10 with epochs range from 100 to 300.
5.2 Classification Performance
5.2.1 Classification Error
@@
The Tab. 2 presents the classification error (%) on the test set for various methods across different datasets and models. FCL consistently achieves the lowest classification error in most cases across all datasets and models, demonstrating its effectiveness in reducing test errors, especially in CIFAR-100 and News datasets.
5.3 Gradient-Based Localization
5.3.1 Grad-CAM Heatmaps on CheXNet
@@
We evaluate the effects of Cross Entropy, Focal Loss, and Focal Calibration Loss on localization performance using Grad-CAM heatmaps, a popular visualization technique for identifying regions of interest in deep learning models (Selvaraju et al., 2017). Fig. 1 presents a comparison of heatmaps generated with each loss function applied to CheXNet (Rajpurkar et al., 2017), a deep learning model for chest X-ray classification. For Cross Entropy, the heatmaps show broader regions of activation, which suggest that the model has less certainty in localizing the anomalies. In contrast, focal loss demonstrates more concentrated regions of activation. This suggests that focal loss sharpens the model’s focus on challenging examples. However, while the activations are more focused, they may not always be aligned with precise boundaries, which could affect the confidence in anomaly localization. Focal Calibration Loss, on the other hand, provides the most well-defined and precise activation areas. The heatmaps generated show a balanced focus, particularly in difficult cases, while also improving probability calibration. This results in clearer boundaries and more accurate localization of anomalies lead to more confident and reliable predictions, making it especially suited for medical diagnosis tasks where precise localization and confidence in predictions are critical. Thus, FCL not only enhances model calibration but also refines anomaly localization, offering substantial benefits for applications in medical imaging where accurate and confident diagnoses are of utmost importance. CheXNet result shows in Appendix A.16.
5.4 Robustness on OoD Data Shift
@@
We assess the robustness of models trained on in-distribution CIFAR-10 and evaluated on two distinct out-of-distribution (OoD) datasets: SVHN and CIFAR-10-C. Tab. 4 presents the performance of Area Under the Receiver Operating Characteristic curve (AUROC) under data shifts. Across most datasets, FCL consistently achieves the highest AUROC, outperforming other models in handling OoD data. This indicates that FCL enhances the model’s ability to distinguish between in-distribution and out-of-distribution samples effectively. This suggests that the FCL generalize well across different types of OoD shifts.
@@
Table 4: Robustness on OoD Dataset Shift. $\uparrow$ AUROC (%) for shifting CIFAR-10 (in-distribution) to SVHN and CIFAR-10-C (OoD).
@@
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
@@
5.5 Discussion
Figure 3: Normalized test loss across training epochs for different loss functions on CIFAR-10 using ResNet-110. The comparison includes Cross Entropy, Focal Loss ($\gamma$=3.0), Dual Focal Loss ( $\gamma$=2.0), and Focal Calibration Loss ($\gamma$=3.0). Focal Calibration Loss with faster convergence and stability.
5.5.1 Faster convergence
@@
5.5.2 Balance between $\gamma$ and $\lambda$
@@
We compared the FCL and FL with different $\lambda$ values while keeping $\gamma=3.0$ constant by metrics including NLL, ECE, AdaECE, and Classwise-ECE both before and after temperature scaling as well as test error for each $\lambda$ in Fig. 6. NLL decreases initially as $\lambda$ increases from 0 to 0.5 but shows fluctuations and a slight increase beyond $\lambda=1.5$. For post-temperature scaling, NLL values are generally lower, indicating improved calibration. ECE decreases significantly when $\lambda$ is 0.5, indicating improved calibration. However, calibration error increases for higher $\lambda$ values, suggesting that too much regularization might degrade calibration performance as depicted in Fig. 6. The experiment shows that $\lambda=0.5$ provides the best calibration performance across all ECE metrics before and after temperature scaling. Adjusting $\gamma$ influences how much the model concentrates on hard examples—higher $\gamma$ increases this focus, which can improve calibration but may also lead to overfitting if excessive. Modifying $\lambda$ controls the weight of the calibration loss component; increasing $\lambda$ emphasizes probability calibration by penalizing deviations between predicted and true probabilities, but overly large values can reduce classification accuracy. Therefore, appropriate tuning of these hyperparameters is crucial to achieving a balance between calibration and classification. We provide table of $\gamma$ and $\lambda$ selection with best calibration error performance in Appendix A.10 and an ablation study of L1/L2 norm in Appendix A.15.
5.5.3 Post-processing gap Reduction
@@
To empirically validate the theoretical reduction in post-processing gap as Theorem 4.5, we compare the ECE/smCE/AdaECE/Classwise-ECE of models trained with focal loss and FCL before and after temperature scaling. As shown in Tab. 4, Tab. 5, Tab. 6 and Tab.7, the focal loss model exhibits a significant reduction in ECE after temperature scaling, indicating a substantial post-processing gap. In contrast, the FCL model shows minimal change in calibration error after temperature scaling closing with 1.0, confirming that FCL reduces the post-processing gap.
@@
Figure 4: Different hyperparameters trained on CIFAR-10 with ResNet-110. The plots show the performance of the FCL with varying calibration parameter $\lambda$ ($\gamma=3.0$) across different metrics: NLL, ECE, AdaECE, and Classwise-ECE. Each plot includes before (blue solid lines) and after (orange dashed lines) temperature scaling, alongside the Test Error (red dotted lines)  [^6] Our FCL reduces to the standard focal loss when $\lambda$ = 0..
6 Related Work
@@
Improving the calibration of deep neural networks (DNNs) is critical for reliable uncertainty estimation. A common strategy involves post-hoc calibration methods, which adjust a pre-trained model’s output probabilities without altering its architecture or training procedure. Platt scaling (Platt et al., 1999), originally designed for binary classifiers, was extended to multi-class settings by applying a linear transformation to logits to improve calibration. More recent methods like Temperature Scaling (Guo et al., 2017) introduce a temperature parameter to adjust the predicted logits, effectively reducing the Expected Calibration Error (ECE) across various models. Histogram Binning (Zadrozny & Elkan, 2002) discretizes the probability space into bins to align confidence scores with empirical frequencies. While effective, these methods require held-out validation data and assume the test distribution matches the calibration data.
6.1 Calibration Strategies
@@
Regularization-based approaches aim to improve calibration during training. Label Smoothing (Müller et al., 2019; Szegedy et al., 2016) mitigates overconfidence by blending true labels with a uniform distribution, softening hard targets to enhance generalization and uncertainty estimation. Additionally, auxiliary losses like the Maximum Mean Calibration Error (MMCE) (Kumar et al., 2018) explicitly minimize the difference between predicted and actual confidence levels. Bayesian techniques offer a principled way to quantify uncertainty. For instance, Deep Ensembles (Lakshminarayanan et al., 2017) aggregate predictions from multiple independently trained models to yield well-calibrated confidence estimates, though they are computationally intensive and less practical in resource-limited settings. Focal Loss (Lin et al., 2017), designed to handle class imbalance in object detection, also affects calibration by concentrating on hard-to-classify examples, which can reduce overconfidence but sometimes result in underconfident predictions (Charoenphakdee et al., 2021). Dual Focal Loss (DFL) (Tao et al., 2023) improves upon this by maximizing the margin between the true class logit and the next highest logit, balancing overconfidence and underconfidence to enhance calibration. Recently, smoothing-based methods have been introduced to provide continuous calibration measures. The Smooth Calibration Error (smCE) (Błasiok & Nakkiran, 2023) uses kernel-based smoothing to create a differentiable measure of calibration error, overcoming the limitations of discrete binning in traditional ECE calculations.
6.2 Calibration in Transformers
Calibration for transformer-based architectures, now prevalent in both vision and language domains, is increasingly essential for reliable deployment in high-stakes applications. In NLP, foundational transformer models such as BERT (Devlin, 2018), GPT (Brown, 2020), and T5 (Raffel et al., 2020) often lead the field in accuracy yet suffer from miscalibrated confidence (Desai & Durrett, 2020; Jiang et al., 2020), especially under out-of-distribution conditions (Mielke et al., 2022). Likewise, vision transformers (ViTs) exhibit calibration issues under distribution shifts (Ovadia et al., 2019), adversarial attacks (Shao et al., 2021), and low-data scenarios (Minderer et al., 2021). These limitations underscore the need for robust calibration methods to enhance reliability in domains such as autonomous driving (Feng et al., 2021) and medical imaging (Mehrtash et al., 2020).
7 Conclusions